Mandy begins bicycling west at 30 mi/h at 11 a.m. If Liz leaves the same point 20 minutes later bicycling west at 36 mi/h, at what time will Liz catch up with Mandy?
Answer by checkley77(12844)

D=RT
D=30T
D=36(T-1/3)
30T=36(T-1/3)
30T=36T-12
30T-36T=-12
-6T=-12
T=-12/-6
T=2 HOURS THE SECOND CYCLIST WILL CATCH THE FIRST CYCLIST.
PROOF:
30*2=36(2-1/3)
60=36*5/3
60=12*5
60=60